Bond Amortization = Bond Discount or Premium / Number of Interest Periods Interest Paid = Face...

  • Bond Amortization = Bond Discount or Premium / Number of Interest Periods

  • Interest Paid = Face Amount of Bonds x Stated Interest Rate

  • Interest Expense = Interest Paid + Discount ( or – Premium) Amortization

On October 1, 2018 ABC issued 5%, 10-year bonds with a face value of $4,000,000 at 104. Interest is paid on October 1 and April 1, with any premiums or discounts amortized on a straight-line basis. What is interest expense for 2018? Assume ABC Company uses straight-line amortization.

N =  

Bond premium =

Bond Amortization = Bond Premium / Number of Interest Periods =  

Interest Paid = Face Amount of Bonds x Stated Interest Rate =  

Interest Expense = Interest Paid - premium amortization =  
Answer #1
N = 20
Bond Premium = ($4000000*104/100) -40000000
=$160000
Bond Amortization = Bond Premium / Number of Interest Periods =$160000/20 = $8000
Interest Paid = Face Amount of Bonds x Stated Interest Rate = $4000000*5%/2 =$100000
Interest Expense = Interest Paid - premium amortization = $100000-8000 =$92000
Answer #2

SOLUTION :


Face value of the bond = 4000000


Premium amount = 4000000 * (104 - 100)/100 = 160000


Maturity period of the bond = 10 years.


Coupon rate is 5% semi-annually.


Hence, coupon amount per period of payments

= Interest expense on the bond per period

= 4000000 * 0.05/2 

= 100000


Premium of 160000 is to be amortised for 20 semi-annual periods


As per straight line method amortisation per period

= 160000 / 20 

= 8000


So, total expense (considered as interest) per period

= Coupon amount - amortised premium

= 100000 -  8000

= 92000.



Bond  issued on October 1, 2018, so in the year 2018 coupon is paid only once on Oct. 1, 2018 itself.


Hence interest expense in 2018 :


= Coupon amount - Amortised premium

= 100000 - 8000 

= 92000 (ANSWER). 

answered by: Tulsiram Garg